OpenStudy (anonymous):

The function f(x) = 6x is an exponential function with base 6, f(-2)=?

OpenStudy (kinggeorge):

is that \(f(x) = 6^x\)? Or \(f(x)=6x\)?

OpenStudy (anonymous):

f(x)=6x

OpenStudy (anonymous):

wait 6^x

OpenStudy (kinggeorge):

Well, if you have a negative exponent, then\[x^{-a}={1\over{x^a}}\]so\[f(-2)=6^{-2} = {1\over{6^2}}\]Could you simplify that the rest of the way?

OpenStudy (anonymous):

thats what i get too but it keeps saying i am wrong

OpenStudy (kinggeorge):

You might have to convert to a decimal. In which case it would be .0277777777777777...

OpenStudy (anonymous):

yep did that too. well i appreciate the help

OpenStudy (kinggeorge):

Talk to your teacher then.